### v1

@@ 2.4 Flux temperature
The temperature in these growths was controlled by changing the applied power of the laser system. The steady-state power was generally set relative to the power level needed to completely melt the flux pellet, which varied directly with the mass. In early trials using a FeCr flux and a h-BN feed rod, the laser power during steady-state growth was set substantially higher than the initial laser power needed to melt the flux (typically an $\approx$ 50% increase). In later growths, the laser power during steady state growth was set to a lower value, typically $\approx$ 5 – 10% over the initial melting power. This lower laser power was preferable, as it was still sufficient for h-BN dissolution by the flux and led to less volatility. The ability to tilt the laser diodes in the PARADIM furnace allows for adjustment of the temperature gradient across the interface between the molten flux and the upper rod. A small inclination ($\phi$ = 2°) enhanced dissolution of the feed rod in the TSFZ geometry and was therefore employed in some TSFZ growths. All LPG growths were conducted with no inclination ($\phi$ = 0°).
@@ 3.1 Micro computed X-ray tomography
Micro computed X-ray tomography (microCT) data were collected using a Bruker SkyScan 1172 operating at 100 kV and 100 $\mu$A with downstream Al and Cu filters in place. The pixel size ranged from 2 – 6 $\mu$m, depending on the highest level of magnification which kept the sample in the field of view. Data was collected by rotating the sample 360° and collecting exposures every 0.25°– 0.35°. Reconstruction of the two-dimensional exposures into a full three-dimensional data set was done using the NRecon software [58] with ring artifact removal and beam hardening correction algorithms applied. The reconstructed data consists of a collection of voxels, each containing an absorption value. This data was then visualized using the Dragonfly software, Version 2021.1 for Windows (Object Research Systems Inc, Montreal, Canada, 2020; software available at http://www.theobjects.com/dragonfly).
@@ 3.2 Raman spectroscopy
Raman scattering experiments were conducted at room temperature ($T$ = 295 K) in the backscattering geometry. Most measurements were conducted using a Horiba LabRAM HR Evolution Raman microscope system with a $\lambda$ = 532 nm laser. The laser light was focused by a 100$\times$ objective lens to a spot size of $\approx$ 1 $\mu$m and the laser power was kept below 0.6 mW to avoid laser-induced heating. The scattered light was dispersed by an 1800 grooves/mm grating and detected by a thermoelectrically cooled CCD. Additional measurements (Fig. S7) were conducted using a Horiba JY T64000 spectrometer equipped with an Olympus microscope and a $\lambda$ = 514.6 nm laser. The light was focused to a spot size of $\approx$ 2 $\mu$m and the laser power was kept below 1 mW. In some cases samples were cut and polished prior to measurement, while in other cases the samples were measured as-synthesized; this distinction is noted in the relevant figure captions.
@@ 3.3 Photoluminescence spectroscopy
Photoluminescence (PL) spectroscopy in the deep UV range was performed at low temperature ($T$ = 8 K) with the samples held on the cold finger of a closed-cycle cryostat. The PL was excited above the bandgap by the fourth harmonic of a CW mode-locked Ti:sapphire oscillator with 140 fs-pulses at 80 MHz repetition rate. The detection system was composed of a $f$ = 500 mm Czerny–Turner monochromator, equipped with an 1800 grooves/mm grating blazed at 250 nm, and a CCD camera (Andor Newton 920). The spectra were recorded across the energy range 3.8 – 6 eV with an excitation energy of 6.2 eV ($\lambda$ = 198 nm) and an incident power of around 40 $\mu$W. All samples were cut and polished before measurement.
@@ 3.5 Scanning electron microscopy
Scanning electron microscope (SEM) backscattered-electron images were collected using a JEOL JSM-IT100 SEM, with a 20 kV acceleration voltage. Samples (either cut and polished or unmodified) were mounted in the SEM using carbon tape on an Al stud. The crystal composition was measured by energy dispersive X-ray spectroscopy (EDS)– with the same acceleration voltage – to collect both spot scans and two-dimensional maps (Fig. S3).
@@ 3.6 Magnetometry
Magnetization versus field data (Fig. S5) were collected at $T$ = 300 K across the range $\mu_{0}H$ = ±7 T using a superconducting quantum interference device (SQUID) magnetometer (Quantum Design, MPMS3) operating in the DC mode. Samples were mounted either by wrapping in thin polypropylene foil and securing inside a plastic straw ("chunk" sample) or by mounting to a quartz paddle with GE varnish ("flakes" sample).
@@ 4 Results
Figure 2: Optical microscopy images (left) and corresponding microCT data (right) for three representative samples produced in this work. The growth conditions for each sample are provided in the caption. (a) Fe flux, $P$ = 5 bar, 0.1 mm/h, h-BN feed, TSFZ, $\phi$ = 2°, $D_{U}$ = $D_{L}$ = 3 mm. (b) FeCrB0.2 flux, $P$ = 50 bar, 0.4 mm/h, h-BN feed, TSFZ, $\phi$ = 0°, $D_{U}$ = $D_{L}$ = 3 mm. (c) FeCrB0.2 flux, $P$ = 5 bar, 0.1 mm/h, h-BN feed, TSFZ, $\phi$ = 2°, $D_{U}$ = $D_{L}$ = 3 mm. For (a) and (c), the sample broke into two sections upon removal from the furnace. Black arrows in the microCT data indicate the location of the orthogonal slices shown in Fig. 3.
In Fig. 2 we show optical images of samples grown under different conditions side-by-side with corresponding microCT data. Images of other exemplar samples are shown in Fig. S1. The range in morphology and outward appearance of the samples is readily apparent. Different regions of individual samples also varied, such as in the sample shown in Fig 2b, where the surface morphology shows a clear crossover between the upper and lower regions. However, most of the samples show at least some colorless/transparent regions on their outer surface, and all show such regions internally (see optical images shown in Fig. 4, 5 and 6). MicroCT provides an efficient first-pass characterization of these samples, yielding a three-dimensional view of the sample volume that provides insight into the nature and morphology of the crystallized material. The microCT images of the three samples shown in Fig. 2 represent a slice through the center of the samples along the growth direction. In Fig. 3 we show cross-sectional slices in the direction orthogonal to the growth direction, with each slice taken at the corresponding black arrow in Fig. 2. Distinct regions of high and low absorption were present in all samples.
@@
EDS analysis provides elemental information regarding the distinct regions seen in the microCT data. Fig. 4a shows an optical image and corresponding microCT cross-section of an early growth attempt with a FeCr flux. The colorless crystalline regions of the optical image correspond to the low absorption regions seen in the microCT cross-section. The darker matrix surrounding these colorless regions corresponds to the high absorption regions. An SEM back-scattered electron image of the sample is shown in Fig. S3a. EDS spot scans taken within the two regions (Fig. 4b) show distinct compositions: the spectra in the darker matrix region (spot 1) is dominated by emission lines from Fe and Cr, while the spectra in the transparent crystalline region shows clear B and N emission lines, with very minor Fe and Cr contributions. Due to their low atomic numbers we do not attempt a quantitative analysis of the B and N stoichiometry by EDS. In Fig. S6 we show SEM/EDS data on a colorless, crystalline flake extracted from the interior of a sample grown with a FeB0.2 flux (Fig. S1d). EDS spot scans from different regions of this sample show strong B and N emission lines, with no sign of Fe emission.
Figure 4: (a) Optical image of a cut and polished sample from an early growth using FeCr flux, $P$ = 1 bar, 0.1 mm/h, h-BN feed, TSFZ, $\phi$ = 0°, and $D_{U}$ = $D_{L}$ = 6 mm. The microCT image corresponds to a region near to the surface shown in the optical image, though from slightly deeper into the sample volume. (b) EDS spectra collected at the labeled positions in (a). Spot 1 is on the darker matrix and spot 2 is on the colorless crystalline material.
@@
Figure 5: Raman data for three representative h-BN samples showing the E2g intralayer (top) and interlayer (bottom) modes. Optical images of the measured samples are shown as insets. The dashed red line is a fit with a Lorentzian peak on a contstant plus linear background; $x_{0}$ is the fitted position of the Lorentzian. The numbers at the top of each panel (1, 2 and 3) match the labeling in the legend of Fig. 6a. Data sets with the same numerical label come either from the same growth (1 and 3) or from analogous growths (2). Sample 2 was taken from the growth shown in Fig. S1e. A more limited Raman data set collected on a sample from the same growth as sample 2 in Fig. 6 can be found in Fig. S7. Samples 2 and 3 were cut and polished before measurement. Sample 1 was extracted from Stage 2 of the sample shown in Fig. 2a (area enclosed in red line) following soaking in aqua regia and washing with ethanol; no cutting and polishing was done. (a) Intralayer averages: FWHM = 7.7(2) cm-1, $x_{0}$ = 1365.46(4) cm-1. Interlayer averages: FWHM = 1.1(1) cm-1, $x_{0}$ = 51.78(9) cm-1. The atomic motions corresponding to each mode are also shown as the insets. (b) Intralayer averages: FWHM = 8.5(3) cm-1, $x_{0}$ = 1365.49(3) cm-1. Interlayer averages: FWHM = 1.4(2) cm-1, $x_{0}$ = 51.81(3) cm-1. (c) Intralayer averages: FWHM = 11.2(5) cm-1, $x_{0}$ = 1365.36(8) cm-1. Interlayer averages: N/A.
@@
We now turn to spectroscopic characterization of select samples. Raman spectra for three samples grown under different conditions are shown in Fig. 5. We refer to the samples by their numerical label (1, 2 or 3) which match them to the corresponding PL data set in Fig. 6. Comparison of samples 1 and 2 illustrates the effect of changing flux composition (Fe and FeCrB0.2, respectively). Both samples show the intralayer (high-energy, E2g) and interlayer (low-energy, E2g) Raman modes characteristic of h-BN [59], further confirming the crystalline structure. However, while there is no significant difference in peak positions, the full-width-at-half-max (FWHM) for both modes is narrower in sample 1 than in sample 2. Sample 3 shows the broadest intralayer mode, and the interlayer mode is not observed. Additional Raman data exploring the impact of flux inclusions on the Raman spectra are shown in Fig. S8.
@@
Figure 6: (a) PL spectra collected on the three samples shown in Fig. 2 following cutting and polishing. The data are referred to in the main text by the numbers in the legend, with 1 corresponding to the sample in Fig. 2a, 2 to Fig. 2c and 3 to Fig. 2b. Optical images of the measured cut and polished surfaces are shown in the corresponding insets. (b) Comparison of the PL spectrum of the highest-quality TSFZ-grown sample (1) to two different samples grown by the CBF method at Kansas State University.
@@
In Fig. 6 we show normalized PL data for the three samples shown in Fig. 2. The data in the range $E\approx$ 4.8 – 5.7 eV are qualitatively similar for all three samples, showing the same features. These features in the PL spectrum of h-BN have been assigned to bound exciton luminescence due to stacking faults and donor-acceptor pair (DAP) recombination driven by charged defects [60, 61]. In contrast, distinct differences are observed at higher ($E>$ 5.7 eV) and lower ($E<$ 4.8 eV) energies. At lower energy (left lower inset, Fig. 6a), samples 2 and 3 show two clear features centered at $E\approx$ 3.9 eV and 4.1 eV, which have been associated with a carbon-related defect [20, 60, 61, 62, 63], most likely a C2 dimer [65]. Sample 1 shows a substantial reduction in these peaks, with no observable intensity at 3.9 eV. A very slight peak is resolvable at 4.1 eV the height of which, relative to the background on the high energy side, is approximately 14 times smaller than the same feature in sample 2 and 51 times smaller than for sample 3. We also see substantial differences in the higher energy region. Sample 1 clearly shows the sharp peaks associated with phonon-assisted free exciton luminescence [4]. Sample 2 shows resolvable remnants of these features although they are substantially broadened. The signal corresponding signal in sample 3 shows no resolvable peaks.

### v2

@@ 2.4 Flux temperature
The temperature in these growths was controlled by changing the applied power of the laser system. The steady-state power was set relative to the power level needed to completely melt the flux pellet, which varied directly with the mass. In initial experiments using a FeCr flux and a h-BN feed rod, the laser power during steady-state growth was set substantially higher than the initial laser power needed to melt the flux (typically an $\approx$ 50% increase). In later growths, the laser power during steady state growth was set to a lower value, typically $\approx$ 5 – 10% over the initial melting power. This lower laser power was preferable, as it was still sufficient for h-BN dissolution by the flux but led to less volatility. This volatility occurs regardless of flux composition and most likely stems from the finite vapor pressure of Fe and Cr since SEM/EDS measurements (not shown) indicate that it contains the metals utilized in the flux. Anecdotally, the rate of volatilization was observed to be greater (for equal laser power) in the presence of Cr, consistent with the higher vapor pressure of Cr. The ability to tilt the laser diodes in the PARADIM furnace allows for adjustment of the temperature gradient across the interface between the molten flux and the upper rod. A small inclination ($\phi$ = 2°) enhanced dissolution of the feed rod in the TSFZ geometry and was therefore employed in TSFZ growths with h-BN feed rods. In the case of B feed rods, no inclination was used ($\phi$ = 0°) to avoid direct heating of the B crystal and subsequent thermally induced cracking. All LPG growths were conducted with no inclination ($\phi$ = 0°).
@@ 3.1 Micro computed X-ray tomography
Micro computed X-ray tomography (microCT) data were collected using a Bruker SkyScan 1172 operating at 100 kV and 100 $\mu$A with downstream Al and Cu filters in place. MicroCT allows for visualization of the internal structure of the grown samples, providing a straightforward means of evaluating flux incorporation and crystal morphology to illuminate the effects of different growth parameters. The pixel size ranged from 2 – 6 $\mu$m, depending on the highest level of magnification which kept the sample in the field of view. Data was collected by rotating the sample 360° and collecting exposures every 0.25°– 0.35°. Reconstruction of the two-dimensional exposures into a full three-dimensional data set was done using the NRecon software [58] with ring artifact removal and beam hardening correction algorithms applied. The reconstructed data consists of a collection of voxels, each containing an absorption value. This data was then visualized using the Dragonfly software, Version 2021.1 for Windows (Object Research Systems Inc, Montreal, Canada, 2020; software available at http://www.theobjects.com/dragonfly).
@@ 3.2 Raman spectroscopy
Raman spectroscopy data was collected in order to evaluate the crystallinity of grown crystals via the metric of the full-width-at-half-max (FWHM) of the measured phonon modes. All measurements were conducted at room temperature ($T$ = 295 K) in the backscattering geometry. Most measurements were conducted using a Horiba LabRAM HR Evolution Raman microscope system with a $\lambda$ = 532 nm laser. The laser light was focused by a 100$\times$ objective lens to a spot size of $\approx$ 1 $\mu$m and the laser power was kept below 0.6 mW to avoid laser-induced heating. The scattered light was dispersed by an 1800 grooves/mm grating and detected by a thermoelectrically cooled CCD. Additional measurements (Fig. S9) were conducted using a Horiba JY T64000 spectrometer equipped with an Olympus microscope and a $\lambda$ = 514.6 nm laser. The light was focused to a spot size of $\approx$ 2 $\mu$m and the laser power was kept below 1 mW. In some cases samples were cut and polished prior to measurement, while in other cases the samples were measured as-synthesized; this distinction is noted in the relevant figure captions.
@@ 3.3 Photoluminescence spectroscopy
Photoluminescence (PL) spectroscopy in the deep UV range was performed at low temperature ($T$ = 8 K) with the samples held on the cold finger of a closed-cycle cryostat. In addition to the intrinsic PL signal, various extrinsic (i.e. disorder-driven) PL processes are known to occur in h-BN. Evaluation of the corresponding PL signals provides a more detailed probe of crystal quality. The PL was excited above the bandgap by the fourth harmonic of a CW mode-locked Ti:sapphire oscillator with 140 fs-pulses at 80 MHz repetition rate. The detection system was composed of a $f$ = 500 mm Czerny–Turner monochromator, equipped with an 1800 grooves/mm grating blazed at 250 nm, and a CCD camera (Andor Newton 920). The spectra were recorded across the energy range 3.8 – 6 eV with an excitation energy of 6.2 eV ($\lambda$ = 198 nm) and an incident power of around 40 $\mu$W. Unless stated otherwise, all samples were cut and polished before measurement.
@@ 3.5 Scanning electron microscopy
Scanning electron microscope (SEM) backscattered-electron images were collected using a JEOL JSM-IT100 SEM, with a 20 kV acceleration voltage. Samples (either cut and polished or unmodified) were mounted in the SEM using carbon tape on an Al stud. The crystal composition was measured by energy dispersive X-ray spectroscopy (EDS) – with the same acceleration voltage – to collect both spot scans and two-dimensional maps.
@@ 3.6 Magnetometry
Taking advantage of the large difference in the magnetic response between h-BN and the various flux compositions, magnetization versus field data were collected as an indirect means of further investigating the nature of flux incorporation. Data were taken at $T$ = 300 K across the field range $\mu_{0}H$ = ±7 T using a superconducting quantum interference device (SQUID) magnetometer (Quantum Design, MPMS3) operating in the DC mode. Samples were mounted either by wrapping in thin polypropylene foil and securing inside a plastic straw ("chunk" sample) or by mounting to a quartz paddle with GE varnish ("flakes" sample).
@@ 4 Results
Figure 2: Optical microscopy images (left) and corresponding microCT data (right) for three representative samples produced in this work. The growth conditions for each sample are provided in the caption. The dashed black arrow on the left shows the growth direction for all samples. The solid black arrows on the microCT images indicate the location of the orthogonal slices shown in Fig. 3, with higher numbers corresponding to regions from later in the growth. The color scale is schematic and applies to all of the presented microCT data. (a) Fe flux, $P$ = 5 bar, vgrow = 0.1 mm/h, h-BN feed, TSFZ, $\phi$ = 2°, $D_{U}$ = $D_{L}$ = 3 mm. (b) FeCrB0.2 flux, $P$ = 5 bar, vgrow = 0.1 mm/h, h-BN feed, TSFZ, $\phi$ = 2°, $D_{U}$ = $D_{L}$ = 3 mm. (c) FeCrB0.2 flux, $P$ = 50 bar, vgrow = 0.4 mm/h, h-BN feed, TSFZ, $\phi$ = 0°, $D_{U}$ = $D_{L}$ = 3 mm. For (a) and (b), the sample broke into two sections upon removal from the furnace. In both cases section 1 is from earlier in the growth and section 2 is from later in the growth.
In Fig. 2 we show optical images of samples (1, 2 and 3) grown under different conditions side-by-side with corresponding microCT data. Optical images and corresponding microCT data of other exemplar samples are shown in Fig. S1 and Fig. S2, respectively. The range in morphology and outward appearance of the samples is readily apparent. Different regions of individual samples also varied, such as in the sample shown in Fig 2c, where the surface morphology shows a clear crossover between the upper and lower regions. However, most of the samples show at least some colorless/transparent regions on their outer surface, and all show such regions internally (see optical images shown in Fig. 4, 5 and 6). MicroCT provides an efficient first-pass characterization of these samples, yielding a three-dimensional view of the sample volume that provides insight into the nature and morphology of the crystallized material. The microCT images of the three samples shown in Fig. 2 represent a slice through the center of the samples along the growth direction. In Fig. 3 we show cross-sectional slices in the direction orthogonal to the growth direction, with each slice taken at the corresponding black arrow in Fig. 2. Distinct regions of high and low absorption were present in all samples.
@@
EDS analysis provides elemental information regarding the distinct regions seen in the microCT data. Fig. 4a shows an optical image and corresponding microCT cross-section for a sample grown from an FeCr flux (sample 4, not shown in Fig. 2 or 3). The colorless crystalline regions of the optical image correspond to the low absorption regions seen in the microCT cross-section. The darker matrix surrounding these colorless regions corresponds to the high absorption regions. An SEM back-scattered electron image of the sample is shown in the SI, in Fig. S3a. EDS spot scans taken within the two regions (Fig. 4b) show distinct compositions: the spectra in the darker matrix region (spot 1) is dominated by emission lines from Fe and Cr, while the spectra in the transparent crystalline region shows clear B and N emission lines, with very minor Fe and Cr contributions. Due to their low atomic numbers we do not attempt a quantitative analysis of the B and N stoichiometry by EDS.
Figure 4: (a) Optical image of a cut and polished section from sample 4, grown by TSFZ using an FeCr flux, $P$ = 1 bar, vgrow = 0.1 mm/h, h-BN feed, $\phi$ = 0°, and $D_{U}$ = $D_{L}$ = 6 mm. The microCT image corresponds to a region near to the surface shown in the optical image, though from slightly deeper into the sample volume. (b) EDS spectra collected at the labeled positions in (a). Spot 1 is on the darker matrix and spot 2 is on the colorless crystalline material.
@@
The microCT and SEM data also demonstrate the clear occurrence of macroscopic regions of trapped flux (segregated from the h-BN regions) within the grown samples. The volume fraction of trapped flux is quantifiable via segmentation of the microCT data based on absorption values, as shown and described in Fig. S6. This analysis yields the following flux volume fractions for the samples shown in Fig. 2: sample 1 = 2 volume %, sample 2 = 38 volume %, and sample 3 = 14 volume %. In the case of samples grown at $P$ = 1 – 7 bar with $v_{grow}$ = 0.1 mm/h, flux residue is often observed on the exterior of the sample as a surface layer (grey regions seen in Fig. 2a,b as well as Fig. S1a), as well as internally as inclusions. The amount of exterior flux tends to be reduced at the higher end of this pressure regime. For samples grown at $P$ = 50 bar with $v_{grow}$ = 0.4 mm/h, the trapped flux appears exclusively in the internal volume of the sample, with the exterior surface totally covered in h-BN. This behavior is shown in Fig. 2c , and was also observed in other samples (not shown) grown under the same conditions. As a comparison to the bulk samples, in Fig. S5 we show SEM/EDS data (as well as other characterization) on a transparent dendritic flake extracted from a sample grown with an FeB0.2 flux. EDS spot scans from different regions of this sample show strong B and N emission lines, with no sign of Fe emission, indicating that minimal Fe enters the h-BN lattice itself during TSFZ growth. Additional investigation of flux incorporation into the grown sample by magnetization measurements is shown in the SI (Fig. S7).
@@
Figure 5: Raman data for three representative TSFZ-grown h-BN samples showing the E2g intralayer (top) and interlayer (bottom) modes. The dashed red line is a fit with a Lorentzian peak on a contstant plus linear background; $x_{0}$ is the fitted position of the Lorentzian. The labels above each set of panels indicates the measured sample (inset optical image) which are those shown in Fig. 2. (a) Intralayer averages: FWHM = 7.7(2) cm-1, $x_{0}$ = 1365.46(4) cm-1. Interlayer averages: FWHM = 1.1(1) cm-1, $x_{0}$ = 51.78(9) cm-1. The measured piece was extracted from section 2 of this sample (area enclosed in red line in Fig. 2a) following soaking in aqua regia and washing with ethanol; no cutting and polishing was done. The ball-and-stick insets show the atomic motions corresponding to each phonon mode. (b) Intralayer averages: FWHM = 8.4(2) cm-1, $x_{0}$ = 1365.63(4) cm-1. Interlayer averages: FWHM = 1.4(2) cm-1, $x_{0}$ = 52.00(2) cm-1. The measured piece is from section 1 of this sample and was cut and polished before measurement. Due to the low intensity, spot 1 was excluded from the interlayer averages. (c) Intralayer averages: FWHM = 11.2(5) cm-1, $x_{0}$ = 1365.36(8) cm-1. Interlayer averages: N/A. The measured piece was cut and polished before measurement.
@@
We now turn to spectroscopic characterization of select samples. Raman spectra for three samples grown under different conditions are shown in Fig. 5. Comparison of samples 1 and 2 illustrates the effect of changing flux composition (Fe or FeCrB0.2, respectively). Both samples show the intralayer (high-energy, E2g) and interlayer (low-energy, E2g) Raman modes characteristic of h-BN [59], further confirming the crystalline structure. However, while there is no significant difference in peak positions, the FWHM for both modes is narrower in sample 1 than in sample 2. Sample 3 shows the broadest intralayer mode, and the interlayer mode is not observed. Additional Raman data exploring the impact of flux inclusions on the Raman spectra are shown in Fig. S8. These data indicate that the Raman spectra of the h-BN grains are not strongly modified near the interface with the trapped flux.
@@
Figure 6: (a) PL spectra collected on three TSFZ-grown samples shown in Fig. 2 and Fig. 3, following cutting and polishing. The numbers in the legend correspond to the sample numbers. The data are normalized to the value of the signal at $E$ = 5.77 eV. The dashed black line shows the energy of the indirect gap ($E_{g}$ = 5.955 eV [4]). Optical images of the measured cut and polished surfaces are shown in the corresponding insets. (b) Comparison of the PL spectrum of the highest-quality TSFZ-grown sample (1) to two different samples grown by the CBF method at Kansas State University.
@@
In Fig. 6 we show representative normalized PL data collected on pieces from the same three samples as for the Raman data (Fig. 5). The data in the energy range $E\approx$ 4.8 – 5.7 eV are qualitatively similar for all three samples, showing the same features. These features in the PL spectrum of h-BN have been assigned to bound exciton luminescence due to stacking faults and donor-acceptor pair (DAP) recombination driven by charged defects [60, 61]. In contrast, distinct differences are observed at higher ($E>$ 5.7 eV) and lower ($E<$ 4.8 eV) energies. At lower energy (left lower inset, Fig. 6a), samples 2 and 3 show two clear peaks centered at $E\approx$ 3.9 eV and 4.1 eV, which have been associated with a carbon-related defect [20, 60, 61, 62, 63], most likely a C2 dimer [64]. Sample 1 shows a substantial reduction in these peaks, with no observable intensity at 3.9 eV. A very slight peak is resolvable at 4.1 eV the height of which, relative to the background on the high energy side, is approximately 14 times smaller than the same feature in sample 2 and 51 times smaller than for sample 3. We also see substantial differences in the higher energy ($E>$ 5.7 eV) region. Sample 1 clearly shows the sharp peaks associated with phonon-assisted free exciton luminescence [4]. Sample 2 shows resolvable remnants of these features although they are substantially broadened. The signal corresponding signal in sample 3 shows no resolvable peaks.